A pebble is dropped from the top of a 25-m tall building. How long will it take before the


pebble hits the ground?

1
Expert's answer
2021-10-31T18:15:20-0400
"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\times25\ m}{9.8\ \dfrac{m}{s^2}}}=2.26\ s."